OpenStudy (anonymous):

Marsha has a bag that contains 4 green marbles, 8 yellow marbles, and 20 red marbles. If she chooses one marble from the bag, what is the probability that the marble is not red?

OpenStudy (anonymous):

@DaBest21

OpenStudy (jadeishere):

Okay, what you need to do is turn it into a fraction add together the amount of marbles 20 + 4 + 8 = 32 20/32 = red 8/32 = yellow 4/32 = green so green and yellow are not red so add together 8/32 + 4/32 = 12/32 then simplify that into 6/16 then simplify it further 3 /8 is the probability it will not be red :)